# arXiv:2603.01270v1[eess.AS] 01 Mar 2026

### v1

@@ 2.3 Demographic Validation
To validate that VoxKnesset captures meaningful speaker variation, we probe the demographic signal encoded in pretrained speech representations. We extract 1024-dimensional WavLM-Large embeddings (mean-pooled) and fit Ridge classifiers for gender, religion (4 classes), and birthplace (6 groups), see table 1. Gender prediction is near-perfect (99.9%). Religion and birth-region accuracies (94.8% and 84.9%) are close to the majority baselines, reflecting natural parliamentary class imbalance; however, per-class analysis reveals a meaningful signal — Muslim speakers are identified with 75.2% recall, may reflect L1 transfer from Arabic, and FSU-origin speakers with 78.4% recall, consistent with detectable Russian accent features. These results suggest that VoxKnesset's demographic metadata may support future research into accent detection and demographic bias auditing in speech systems.

### v2

# Marmor Zulti Krongauz Gabet Snapir Lifshitz Segal

@@ 2.3 Demographic Validation
To validate that VoxKnesset captures meaningful speaker variation, we probe the demographic signal encoded in pretrained speech representations. We extract 1024-dimensional WavLM-Large embeddings (mean-pooled) and fit Ridge classifiers for gender, religion (4 classes), and birthplace (6 groups), see table 1. Gender prediction is near-perfect (99.9%). Religion and birth-region accuracies (94.8% and 84.9%) are close to the majority baselines, reflecting natural parliamentary class imbalance; however, per-class analysis reveals a meaningful signal — Muslim speakers are identified with 75.2% recall, may reflect L1 transfer from Arabic, and former Soviet Union origin speakers with 78.4% recall, consistent with detectable Russian accent features. These results suggest that VoxKnesset's demographic metadata may support future research into accent detection and demographic bias auditing in speech systems.